OpenStudy (anonymous):

can someone explain steps to get from sinL(1 + cosB) to 4sinL/2 cos L/2 cos^2 B/2

OpenStudy (anonymous):

sinl=2sin(l/2)cos(l/2) and cosb=(2cos^2(b/2)-1) so sinl(1+cosb) = 2sin(l/2)cos(l/2)(1+2cos^(b/2)-1) = 2sin(l/2)cos(l/2)(2cos^2(b/2)) = 4sin(l/2)cos(l/2)cos^2(b/2)

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

welcome..

OpenStudy (anonymous):

can you explain me how did you get cosb = cos^2b -1